Find the curvature of the parabolic curve r(t) =<t,t^2>.
1
Expert's answer
2020-09-25T10:27:42-0400

r(t) = t , t^2

r,(t) = 1, 2t

r,,(t) = 0 , 2

r, (t)×r,,(t) = (0,0,2)

|r,(t)×r,,(t)| = (0+0+4)1/2 = 41/2

|r,(t)| = (1+4t2)1/2

|r,(t)|3 = (1+4t2)3/2

Curvature = |r,(t)×r,,(t)|/|r,(t)|3

= [(4)1/2/(1+4t2)3/2] ans.